OpenStudy (zmudz):

Let \((a_k)\) be a sequence of integers such that \(a_1 = 1\) and \(a_{m + n} = a_m + a_n + mn\) for all positive integers \(m\) and \(n\). Find \(a_{12}\).

ganeshie8 (ganeshie8):

Let \(m=1\), \[a_{n+1} = a_n+n+1 \\\implies a_{n+1}-a_n = n+1\\\implies \sum\limits_{n=1}^{11} (a_{n+1}-a_n) = \sum\limits_{n=1}^{11} (n+1)\\ \] left side telescopes and right side can be easily evaluated